What is the pH of 6.0 M aqueous solution of NH3 (ammonia). It's Kb = 1.8 x 10-5.
Answer #1

NH3 + H20 6.0 NHW +OH 0 0 Initial - t=o At tat 6.o-x x x Kb = [NHy t] [OH-] Сонд] Kb= 1.8x10-5 1.8x10-5 = Tuba [x] [2] [6.o-x